OpenStudy (anonymous):

Ami bought some gardening tools. She paid $15 for a rake, $18 for a garden hose, and $47 for a wheelbarrow. What percent of the total amount that Ami paid was the price of the rake? Round to the nearest percent, if necessary. % 5 MEDAL REWARD WHO HELPS AND GETS IT RIGHT

OpenStudy (anonymous):

Any options or Essay?

OpenStudy (anonymous):

So I added all 15 + 18 + 47 = 80 then I did: 80 / 15 % = 12 (12 percent maybe?)

OpenStudy (anonymous):

so are u sure

OpenStudy (anonymous):

yes, but is 12 an option? is this an essay

OpenStudy (amistre64):

this process is almost valid. 15 out of 80 is not 12